Lat f -2, g - with th de fra, hon & find the proation f rto 9 tha intorual Fn Over
Answer #1


projection fo f (x) onto g(x) is , f(x), g(x)), (8 (3):(3)}=(x) (1(3).E(=))= }(2 - x)(2x + 1)ds =int((72 - x)(2x – 1),x,-1,1)